# Chinese Academy of Sciences, Beijing, China

### v2

@@ 4.1 Construction of a one-way function
Theorem 4.5.
There is a total computable random-preserving surjection $f:2^{\omega}\to 2^{\omega}$ such that every randomized inversion of $f$ computes $\emptyset^{\prime}$.
Proof..
Let $f,L$ be as in Theorem 4.4 and let $z$ be a real that computes a randomized inversion $g:\subseteq 2^{\omega}\times 2^{\omega}\to 2^{\omega}$ of $f$. Then $\mu(L)>0$ and
$\big{(}{y\oplus r\in L\ \textrm{ and $y\oplus r$ has density $1$ in $L$}}\big{)}\ \Rightarrow\ \emptyset^{\prime}\leq_{T}y\oplus r\oplus z$
as in the proof of Theorem 4.4. By the Lebesgue density theorem
[…]
@@
Let $f,L$ be as in Theorem 4.4 and
$B_{q}\ :=\ \{\hskip 0.85355pt{y}:{\mu(\{\hskip 0.85355pt{r}:{y\oplus r\in L}\hskip 0.85355pt\})>q}\hskip 0.85355pt\}$
so $\mu(L)=\mu(B_{q})=0$ for each $q>0$. Let $z$ be such that:
$\mu(\{\hskip 0.85355pt{r}:{z\oplus r\in L}\hskip 0.85355pt\})>q>0$
for a rational $q>0$ so $z\in B_{q}$. If $g$ is partial computable then $B_{q}$ is a null $\Pi^{0}_{2}$ class, so $z$ is not weakly 2-random. If $g$ is total computable then $B_{q}$ is a null $\Pi^{0}_{1}$ class, so $z$ is not weakly 1-random. ∎
We do not know if Proposition 4.6 (i) holds for incomplete random $z$.
Question.
Characterize the class of reals where arbitrary computable random-preserving $f:2^{\omega}\to 2^{\omega}$ can be inverted with positive probability.
4.2 Injectivity and one-way functions
We consider the extent to which one-way functions fail to be injective. Our observations hold generally with respect to inversions on one variable.
@@ 4.2 Injectivity and one-way functions
Every (weakly) one-way total computable function is almost nowhere injective.
@@
Let $f:2^{\omega}\to 2^{\omega}$ be computable and $E_{f}:=\{\hskip 0.85355pt{y}:{|{f^{-1}(y)}|=1}\hskip 0.85355pt\}$.
Assuming that $\mu(E_{f})>0$ it suffices to show that $f$ is not weakly one-way. By Theorem 3.1 there exists partial computable $h$ with domain $E_{f}$ and $\forall y\in E_{f},\ f(h(y))=y$. Then $f$ is not weakly one-way. ∎
We show that the inverse images $f^{-1}(y)$ of one-way functions $f$ are perfect uncountable sets almost everywhere in their range. For this proof, note that Theorem 3.1 relativizes to any neighborhood $\llbracket\sigma\rrbracket$.
@@
Given total computable $f:2^{\omega}\to 2^{\omega}$ with $\mu(D)>0$ where
$D:=\{\hskip 0.85355pt{y}:{\textrm{$f^{-1}(y)$ is not perfect}}\hskip 0.85355pt\}$
we show that $f$ is not weakly one-way. Since countable sets of reals are not perfect, this implies the statement.
For each $\sigma$ let $f_{\sigma}$ be the restriction of $f$ to $\llbracket\sigma\rrbracket$ and apply Theorem 3.1 to get a partial computable $g_{\sigma}$ which inverts $f_{\sigma}$ on $E_{\sigma}:=\{\hskip 0.85355pt{y}:{|{f_{\sigma}^{-1}(y)}|=1}\hskip 0.85355pt\}$.
If $y\in D$, then the closed set $f^{-1}(y)$ has an isolated path, so $\exists\sigma,\ y\in E_{\sigma}$. So
$D\subseteq\bigcup_{\sigma}E_{\sigma}$
and since $\mu(D)>0$, there exists $\sigma$ with $\mu(E_{\sigma})>0$. By the choice of $g_{\sigma}$:
$\mu(\{\hskip 0.85355pt{y}:{g_{\sigma}(y)\downarrow\ \wedge\ \ f(g_{\sigma}(y))=y}\hskip 0.85355pt\})>0$
so $f$ is not weakly one-way. ∎
5 Injectivity and hardness of inversion
We have seen in §3 that injective computable $f:2^{\omega}\to 2^{\omega}$ have computable inversion. This fails for partial computable functions.
Theorem 5.1.
Proof..
Let $f:\subseteq 2^{\omega}\to 2^{\omega}$ be defined by $f(x)\simeq p(x)\oplus q(x)$ where
$p(x;\langle n,s\rangle):=$$\begin{cases}x(n)&\textrm{if $n\in\emptyset^{\prime}_{s+1}-\emptyset^{\prime}_{s}$}\\[2.84544pt] 0&\textrm{otherwise}\end{cases}$
$q(x;n):\simeq$$\begin{cases}0&\textrm{if $\forall i\leq n,\ \big{(}{x(i)=0\ \vee\ i\in\emptyset^{\prime}}\big{)}$}\\[2.84544pt] \uparrow&\textrm{otherwise}\end{cases}$
and $\simeq$ denotes equality when both sides are defined. Then
  • $f$ is partial computable
  • if $f(x)\downarrow$, $f(z)\downarrow$ and $x\neq z$ then $x,z$ can only differ on positions in $\emptyset^{\prime}$.
In the latter case $p(x)\neq p(z)$ and $f(x)\neq f(z)$, so $f$ is injective.
Given partial $g:\subseteq 2^{\omega}\to 2^{\omega}$ such that
$f(g(y\oplus z))=y\oplus z\hskip 8.5359pt\textrm{ for all $y\oplus z\in f(2^{\omega})$}$
we show $g\geq_{T}\emptyset^{\prime}$. Since $f(0^{\omega})\downarrow=0^{\omega}$ and $f$ is injective, $g(0^{\omega})\downarrow=0^{\omega}$.
To decide if $n\in\emptyset^{\prime}$ let $u$ be the oracle-use of $g(0^{\omega};n)=0$. We claim that
$n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{u}.$
Assuming otherwise, there is $s>u$ such that $n\in\emptyset^{\prime}_{s+1}-\emptyset^{\prime}_{s}$. Then
$f(0^{n}10^{\omega})\downarrow=y\oplus z\hskip 8.5359pt\textrm{where $y=0^{\langle n,s\rangle}10^{\omega}$ and $z=0^{\omega}$.}$
Since $f$ is injective and $f(g(y\oplus z))=y\oplus z$, we have $g(y\oplus z)=0^{n}10^{\omega}$.
But $y,0^{\omega}$ agree on positions $\langle n,s\rangle>s>u$ so
$10=g(0^{\omega};n)=g(y\oplus z;n)=0^{n}10^{\omega}(n)=$ 1
which is a contradiction. So $g$ computes $\emptyset^{\prime}$. ∎
Inverting a non-injective computable $f:2^{\omega}\to 2^{\omega}$ can be as hard as computing $\emptyset^{\prime}$. This was demonstrated in Theorem 3.3 by a simple definition of a highly non-injective $f$: $f^{-1}(y)$ is uncountable for each $y\in f(2^{\omega})$.
The extent that $f$ fails to be injective can be measured by the distribution of sizes of the $f^{-1}(y)$, $y\in f(2^{\omega})$. We wish to determine the extent that hard-to-invert computable functions $f$ fail to be injective.
@@ 5 Injectivity and hardness of inversion
At the cost of some complexity in $f$ we show:
@@
and the latter holds for the restriction of $f$ in any neighborhood $\llbracket\sigma\rrbracket$.
By Theorem 4.9 every computable two-to-one random-preserving surjection (including $f$ of Theorem 5.3) can be effectively inverted with positive probability. This leaves the possibility that a computable $f$ exists such that:
  1. (i)
  2. (ii)
    no partial computable $g$ inverts $f$ with probability 1.
We give a strong positive answer:
Theorem 5.4.
There is a total computable $f:2^{\omega}\to 2^{\omega}$ such that:
  • $f$ is a two-to-one random-preserving surjection
  • every partial $g\not\geq_{T}\emptyset^{\prime}$ fails to invert $f$ with positive probability
and the latter holds for the restriction of $f$ in any neighborhood $\llbracket\sigma\rrbracket$.
We use the following relativization of Lemma 4.3:
Lemma 5.5.
Let $p$ be a total Turing functional such that $n\mapsto p^{z}(n)$ is injective for each oracle $z$ and define $f,g^{z}:2^{\omega}\to 2^{\omega}$ by
$g^{z}(x;n):=x(p^{z}(n))\hskip 8.5359pt\textrm{and}\hskip 8.5359ptf(x\oplus z):=g^{z}(x)\oplus z.$
Then $f$ is a total computable random-preserving surjection.
@@
Since $p$ is a total Turing functional, $(z,x)\mapsto g^{z}(x)$ and $f$ are computable. By the relativization of Lemma 4.3 to arbitrary oracle $z$:
$g^{z}$ is a $z$-random preserving surjection.
So $f$ is a surjection and if $x\oplus z$ is random, by van Lambalgen’s theorem:
[…]
so $f$ is random-preserving. ∎
5.1 Construction of two-to-one functions
In Theorem 5.3 we define $f(x\oplus z):=h^{z}(x)\oplus z$ where
  • $h^{z}$ selects positions of $x$-bits used in (i.e. copied into) $h^{z}(x)$
  • all but at most one position are used in $h^{z}(x)$.
Input: $x\oplus z$;     Output: $y$ with $f(x\oplus z)=y\oplus z$ 1:   Initialization: $k:=0$, $s:=0$ 2:  while true do 3:     if $k\in\emptyset^{\prime}_{s}$ or $E_{s}^{z}(k)$ then 4:        $y(s):=x(k)$ 5:        $k:=s+1$ 6:     else 7:        $y(s):=x(s+1)$ 8:     end if 9:     $s:=s+1$ 10:  end while
Figure 1: Definition of $f$ given a $z$-computable predicate $E^{z}_{s}(i)$.
Figure 1 depicts $f$ as an algorithm, where $y=h^{z}(x)$. Let $k^{z}_{s}$ be the candidate for the unique unused position at $s$, corresponding to the $k$ in the figure. The update of $k^{z}_{s}$ occurs if:
  • either $k^{z}_{s}\in\emptyset^{\prime}_{s+1}$ which we call a $\emptyset^{\prime}$-permission
  • or $E^{z}_{s+1}(k^{z}_{s})$ which we call a $z$-permission
in both cases candidate $k^{z}_{s}$ is eliminated. Since $k^{z}_{s}$ is non-decreasing:
  1. (i)
    if $\lim_{s}k^{z}_{s}=\infty$ all $x$-positions are used in $h^{z}(x)$
  2. (ii)
    if $\lim_{s}k^{z}_{s}=k_{0}$ all $x$-positions except $k_{0}$ are used in $h^{z}(x)$.
5.2 Proof of Theorem 5.3
We define $f$ according to §5.1 for predicate $z(\langle i,s\rangle)=1$ in place of $E^{z}_{s}(i)$. For each $z$ let $k^{z}_{0}=0$ and
@@ 5.1 Construction of two-to-one functions
is a total computable random-preserving surjection. Also $f$ is two-to-one and
$\forall k,\ |{\{\hskip 0.85355pt{i}:{z(\langle k,i\rangle)=1}\hskip 0.85355pt\}}|=\infty\ \Rightarrow\ \lim_{s}k^{z}_{s}=\infty\ \Rightarrow\ |{f^{-1}(h^{z}(x)\oplus z)}|=1$
as discussed in §5.1. The condition is satisfied for all random $z$, so $f$ is almost everywhere injective and by Theorem 3.1 it has a partial computable almost everywhere inversion.
@@
Then $k_{t}^{z}$ does not get stuck on any number $\neq n$ and
@@
    $k_{s}^{z}$ is updated at $t>s$ iff $n\in\emptyset^{\prime}_{t}-\emptyset^{\prime}_{t-1}$.
To decide if $n\in\emptyset^{\prime}$ we show $n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{\max\{u,s\}}$ where
$g(0^{\omega}\oplus z;2n)\downarrow\hskip 8.5359pt\textrm{with oracle-use $u$.}$
@@
$\forall x,y\ \ \big{(}{f(x\oplus z)=y\oplus z\ \Rightarrow\ y(t)=x(n)}\big{)}$ (7)
due to (b) and the definition of $f$. Let
$y_{0}:=0^{\omega}\hskip 8.5359pt\textrm{and}\hskip 8.5359pty_{1}:=0^{t}10^{\omega}$
so for $i=0,1$, $f(g(y_{i}\oplus z))=y_{i}\oplus z$ and by (7):
@@
which contradicts (8). This concludes the proof of $n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{\max\{u,s\}}$.
Finally we modify the above argument so that $g\geq_{T}\emptyset^{\prime}$ is obtained from the weaker assumption that $g$ inverts $f$ inside a neighborhood $\llbracket\upsilon\oplus\zeta\rrbracket$. Let
@@
Assuming $n>|\zeta|$ the modified $z$ satisfies (a), (b). So the above argument applies to the modified $y_{0},y_{1},u$ and proves $g\geq_{T}\emptyset^{\prime}$ as required.
5.3 Proof of Theorem 5.4
We construct a two-to-one total computable random-preserving surjection $f$ such that any any $g\not\geq_{T}\emptyset^{\prime}$ fails to invert $f$ with positive probability.

### v3

@@ 3.2 Construction of a total one-way surjection
it remains to show that $g$ computes $\emptyset^{\prime}$. Fix $\sigma$ such that
$\mu(L\cap\llbracket\sigma\rrbracket)>\frac{3}{4}\cdot\mu(\llbracket\sigma\rrbracket)$ (4)
which exists by the Lebesgue density theorem [8, Theorem 1.2.3] or by simply taking an open cover $\llbracket V\rrbracket$ of $L$ such that $\mu(\llbracket V\rrbracket-L)<\mu(L)/3$, where $V=\{\sigma_{0},\sigma_{1},\dots\}$ is prefix-free. Then some $\sigma_{i}$ in $V$ must satisfy (4).
We now compute $\emptyset^{\prime}$ using $g$ and the effective enumeration of a set $W$.
To decide if $n\in\emptyset^{\prime}$, we simultaneously compute $g(y,r;2n)$ for all $y,r$:
if $g(y,r;2n)\downarrow$ with oracle-use $u\geq|\sigma|$ enumerate $(y\oplus r)\upharpoonright_{u}$ in $W$.
Then $L\subseteq\llbracket W\rrbracket$ and $W$ is a c.e. prefix-free set with $\forall\tau\in W,\ |\tau|\geq|\sigma|$.
Effectively in $n$ we produce a computable enumeration $(W_{s})$ of $W$ and
  • compute the least $t$ with $\mu(\llbracket W_{t}\rrbracket\cap\llbracket\sigma\rrbracket)>\mu(\llbracket\sigma\rrbracket)/2$
  • compute the length $k$ of the longest string in $W_{t}$
which exist by (4) and $L\subseteq\llbracket W\rrbracket=\bigcup_{t}\llbracket W_{t}\rrbracket$. It remains to show that
$n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{k}.$ (5)
For a contradiction assume that $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$ for some $s>k$.
By the definition of $f$ we have $\forall x,\ f(x;\langle n,s\rangle)=x(2n)$ so
$y\oplus r\in L\implies y(\langle n,s\rangle)=f(g(y\oplus r);\langle n,s\rangle)=g(y\oplus r;2n).$
Fix $\tau\in W_{t}$. Since $\langle n,s\rangle\geq s>k\geq|\tau|$ we have
  • $g(y,r;2n)$ outputs the same result $i$ for $y\oplus r\in\llbracket\tau\rrbracket$
  • only half of the $y\oplus r\in\llbracket\tau\rrbracket$ satisfy $y(\langle n,s\rangle)=i$
so half of $y\oplus r\in\llbracket\tau\rrbracket$ must be outside $L$. Formally:
$\mu(\llbracket\tau\rrbracket-L)\geq\mu(\llbracket\tau\rrbracket)/2$
for each $\tau\in W_{t}$. So $\mu(\llbracket\sigma\rrbracket-L)$ is at least
$\mu(\llbracket W_{t}\rrbracket\cap\llbracket\sigma\rrbracket-L)\geq\mu(\llbracket W_{t}\rrbracket\cap\llbracket\sigma\rrbracket)/2>\mu(\llbracket\sigma\rrbracket)/4$
which contradicts (4), completing the proof of (5). So $g$ computes $\emptyset^{\prime}$.
Since no partial computable $g$ computes $\emptyset^{\prime}$, $f$ is one-way. ∎
one-way functions are not probabilistically invertible on any sufficiently random $z$. We quantify the level of randomness required for this fact.
@@
Let $f,L$ be as in Theorem 3.4 and define
$B_{q}\ :=\ \{\hskip 0.85355pt{y}:{\mu(\{\hskip 0.85355pt{r}:{y\oplus r\in L}\hskip 0.85355pt\})\geq q}\hskip 0.85355pt\}.$
Then $\mu(L)=\mu(B_{q})=0$ for each $q\geq 0$. Let $z$ be such that:
$\mu(\{\hskip 0.85355pt{r}:{z\oplus r\in L}\hskip 0.85355pt\})\geq q>0$
for a rational $q>0$ so $z\in B_{q}$. If $g$ is partial computable then
$y\oplus r\in L\iff\forall n\ \exists s\ f(g(y,r);n)[s]\downarrow=y(n)$
so $L$ is $\Pi^{0}_{2}$. Let $(L_{n,s})$ be a computable family of clopen sets with
$L=\bigcap_{n}\bigcup_{s}L_{n,s}\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\bigcup_{s}L_{n+1,s}\subseteq\bigcup_{s}L_{n,s}$
and $L_{n,s}\subseteq L_{n,s+1}$. Then $B_{q}$ is a null $\Pi^{0}_{2}$ class as it is definable by
$y\in B_{q}\iff\forall n\ \exists s\ \mu(\{r:\{y\oplus r\in L_{n,s}\})\geq q.$
By (2) it follows that $z$ is not weakly 2-random. If $g$ is total computable, then $L$ is a null $\Pi^{0}_{1}$ class. By (2) it follows that $z$ is not weakly 1-random. ∎
3.3 Properties of one-way functions
The proof that the total computable $f$ of §3.2 is one-way relied on the fact that $f$ is not injective. This is not a coincidence: since
  • for each tree $T$ with $[T]=\{x\}$ we have $x\leq_{T}T$ uniformly in $T$
  • if $f$ is a total computable injection then $f^{-1}(y)$ consists of the unique path through a tree which is uniformly computable in $y$
total computable injections have computable inverses. More generally:
Theorem 3.6.
Proof..
Let $\hat{f}$ be a computable representation of $f$. For $y\in E$ the tree
$T_{y}=\{\sigma:\hat{f}(\sigma)\prec y\}$
is uniformly computable in $y$: the map $y\mapsto T_{y}$ is computable. Also
  • $[T_{y}]=f^{-1}(y)$ because $f$ is total
  • if $y\in E$ then $f^{-1}(y)$ is a singleton so $T_{y}$ has a unique path.
The path of any tree $T$ with $|{[T]}|=1$ is computable uniformly in $T$.
So we can compute the unique $x$ in $f^{-1}(y)$ from $y\in E$ uniformly in $y$. ∎
Corollary 3.7.
Partial computable injections need not have computable inverses.
Theorem 3.8.
Proof..
Let $f:\subseteq 2^{\omega}\to 2^{\omega}$ be defined by $f(x):=p(x)\oplus q(x)$ where
$p(x;\langle n,s\rangle):=$$\begin{cases}x(n)&\textrm{if $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$}\\[2.84544pt] \ \ 0&\textrm{otherwise}\end{cases}$
$q(x;n):=$$\begin{cases}0&\textrm{if $\forall i\leq n,\ \big{(}{x(i)=0\ \vee\ i\in\emptyset^{\prime}}\big{)}$}\\[2.84544pt] \uparrow&\textrm{otherwise.}\end{cases}$
Then $f$ is partial computable and
  • if $f(x)\downarrow$, $f(z)\downarrow$, $x\neq z$ then $x,z$ can only differ on positions in $\emptyset^{\prime}$.
In the latter case $p(x)\neq p(z)$ and $f(x)\neq f(z)$, so $f$ is injective.
As in the proof of Proposition 3.2, every inversion $g$ of $f$ computes $\emptyset^{\prime}$. ∎
We now consider the extent to which total computable one-way functions fail to be injective. Our results hold for a weaker type of one-way functions.
@@ 3.3 Properties of one-way functions
Every (weakly) one-way total computable $f:2^{\omega}\to 2^{\omega}$ is almost nowhere injective.
@@
Given $f$ as in the statement suppose that
[…]
By Theorem 3.6 there is a partial computable function $h$ with domain $E_{f}$ and $\forall y\in E_{f},\ f(h(y))=y$. So $f$ is not weakly one-way. ∎
For the following, note that Theorem 3.6 relativizes to any cylinder $\llbracket\sigma\rrbracket$.
@@
Suppose that $f:2^{\omega}\to 2^{\omega}$ is total computable. Since countable sets of reals are not perfect it suffices to show that if
$\mu(D)>0\hskip 8.5359pt\textrm{where}\hskip 5.69046ptD:=\{\hskip 0.85355pt{y}:{\textrm{$f^{-1}(y)$ is not perfect}}\hskip 0.85355pt\}$
then $f$ is not weakly one-way. For each $\sigma$ let
  • $f_{\sigma}$ be the restriction of $f$ to $\llbracket\sigma\rrbracket$ and set $E_{\sigma}:=\{\hskip 0.85355pt{y}:{|{f_{\sigma}^{-1}(y)}|=1}\hskip 0.85355pt\}$
  • $g_{\sigma}$ be partial computable which inverts $f_{\sigma}$ on $E_{\sigma}$
where the existence of the $g_{\sigma}$ follows from Theorem 3.6.
If $y\in D$ the closed set $f^{-1}(y)$ has an isolated path, so
$\exists\rho,\ y\in E_{\rho}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptD\subseteq\bigcup_{\sigma}E_{\sigma}.$
Since $\mu(D)>0$ there is $\sigma$ with $\mu(E_{\sigma})>0$ and by the choice of $g_{\sigma}$:
$\mu(\{\hskip 0.85355pt{y}:{g_{\sigma}(y)\downarrow\ \wedge\ \ f(g_{\sigma}(y))=y}\hskip 0.85355pt\})>0.$
This shows that $f$ is not weakly one-way. ∎
4 Inversions of nearly injective functions
Since one-way injections are well-studied in computational complexity [16, 12] it is interesting to ask if there are one-way injections $f$. Their existence remains unknown but as discussed in §3.3, they cannot be total.
With this motivation, we examine the extent to which non-injectivity is essential in the arguments of §3. We exhibit total computable random-preserving surjections that are hard to invert and are nearly injective.
@@ 4 Inversions of nearly injective functions
By extending the method of §3, in §4.1 we exhibit a total computable two-to-one random-preserving surjection whose inversions compute $\emptyset^{\prime}$ but is nevertheless almost everywhere effectively invertible. In §4.2 we exhibit a two-to-one total computable random-preserving surjection with no effective probabilistic map that inverts it almost everywhere.
4.1 Blueprint for two-to-one functions
Our maps will be of the form $f(x\oplus z):=h^{z}(x)\oplus z=y\oplus z$ where
  • $h^{z}$ selects positions of $x$-bits used in (i.e. copied into) $h^{z}(x)$
  • all but at most one position are used in $h^{z}(x)$.
Input: $x\oplus z$;   Output: $y$ with $f(x\oplus z)=y\oplus z$. 1:   Initialization: $k:=0$, $s:=0$ 2:  while true do 3:     if $k\in\emptyset^{\prime}_{s}$ or $E_{s}^{z}(k)$ then 4:        $y(s):=x(k)$ 5:        $k:=s+1$ 6:     else 7:        $y(s):=x(s+1)$ 8:     end if 9:     $s:=s+1$ 10:  end while
Figure 1: Definition of $f$ given a $z$-computable predicate $E^{z}_{s}(i)$.
The selection of $x$-bits is facilitated by a movable marker $k$ and depends on a computable predicate $E^{z}_{s}(i)$ as shown in Figure 1. Let $k^{z}_{s}$ be the candidate for the unique unused position at $s$, which corresponds to $k$ in Figure 1.
The update of $k^{z}_{s}$ occurs if
  • either $k^{z}_{s}\in\emptyset^{\prime}_{s}$ which we call a $\emptyset^{\prime}$-permission
  • or $E^{z}_{s}(k^{z}_{s})$ which we call a $z$-permission
in which case candidate $k^{z}_{s}$ is eliminated. Since $k^{z}_{s}$ is non-decreasing:
  1. (i)
    if $\lim_{s}k^{z}_{s}=\infty$ all $x$-positions are used in $h^{z}(x)$
  2. (ii)
    if $\lim_{s}k^{z}_{s}=k_{0}$ all $x$-positions except $k_{0}$ are used in $h^{z}(x)$.
We now need a relativization of Lemma 3.3 that applies to this extended form of selective permutation. To this end, we use a fact from [27]:
$x\oplus y$ is random iff $x$ is random and $y$ is $x$-random (6)
also known as van Lambalgen’s theorem.
Lemma 4.2.
Let $p$ be a total Turing functional such that $n\mapsto p^{z}(n)$ is injective for each oracle $z$ and define $f,h^{z}:2^{\omega}\to 2^{\omega}$ by
$h^{z}(x;n):=x(p^{z}(n))\hskip 8.5359pt\textrm{and}\hskip 8.5359ptf(x\oplus z):=h^{z}(x)\oplus z.$
Then $f$ is a total computable random-preserving surjection.
Proof..
Since $p$ is a total Turing functional, $(z,x)\mapsto h^{z}(x)$ and $f$ are computable. By the relativization of Lemma 3.3 to arbitrary oracle $z$ it follows that $g^{z}$ is a $z$-random preserving surjection.
So $f$ is a surjection. If $x\oplus z$ is random, by (6) we get that
[…]
so $h^{z}(x)\oplus z$ is random. Hence $f$ is random-preserving. ∎
We now apply the above framework to prove:
@@ 4.1 Blueprint for two-to-one functions
and the latter holds for the restriction of $f$ in any cylinder $\llbracket\sigma\rrbracket$.
@@
We define $f$ as above with predicate $z(\langle i,s\rangle)=1$ in place of $E^{z}_{s}(i)$.
For each $z$ let $k^{z}_{0}=0$ and
@@
is a total computable random-preserving surjection. Also
$\forall k,\ |{\{\hskip 0.85355pt{i}:{z(\langle k,i\rangle)=1}\hskip 0.85355pt\}}|=\infty\implies\lim_{s}k^{z}_{s}=\infty\implies|{f^{-1}(h^{z}(x)\oplus z)}|=1$
as discussed above, and $f$ is two-to-one. This condition is met for all random $z$, so $f$ is almost everywhere injective. By Theorem 3.6 there is a partial computable almost everywhere inversion of $f$.
@@
Then $k_{t}^{z}$ does not get stuck on any number $\neq n$ in the sense that
@@
    $k_{t}^{z}$ is updated at $t>n$ iff $n\in\emptyset^{\prime}_{t}$.
We show $n\in\emptyset^{\prime}$ iff $n\in\emptyset^{\prime}_{\max\{u,s\}}$, where $u$ is the oracle-use of $g(0^{\omega}\oplus z;2n)$.
@@
$\forall x,y\ \ \big{(}{f(x\oplus z)=y\oplus z\implies y(t)=x(n)}\big{)}$ (9)
due to (b) and the definition of $f$. Let $y_{0}:=0^{\omega}$ and $y_{1}:=0^{t}10^{\omega}$.
Then for $i=0,1$ we have $f(g(y_{i}\oplus z))=y_{i}\oplus z$ and by (9):
@@
which contradicts (10). It follows that $n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{\max\{u,s\}}$.
Finally we modify the above argument so that $g\geq_{T}\emptyset^{\prime}$ is obtained from the weaker assumption that $g$ inverts $f$ inside a cylinder $\llbracket\upsilon\oplus\zeta\rrbracket$. Let
@@
Assuming $n>|\zeta|$ the modified $z$ satisfies (a), (b). So the above argument applies to the modified $y_{0},y_{1},u$ and proves $g\geq_{T}\emptyset^{\prime}$ as required. ∎
4.2 Almost everywhere probabilistic inversions
By Theorem 3.11 every computable two-to-one random-preserving surjection can be effectively inverted with positive probability. This leaves the possibility that a total computable $f$ exists such that:
  1. (i)
  2. (ii)
    no partial computable $g$ inverts $f$ with probability 1.
We construct $f$ with the above properties within the framework of §4.1, starting with the modifications and additional ideas needed to achieve this.